# LtU-ILI: An All-in-One Framework for Implicit Inference in Astrophysics and Cosmology

### v1

@@ 7 Conclusion
In this manuscript, we have explained the theoretical basis of ILI, described the current features of LtU-ILI in detail, and provided benchmarks and science examples on a variety of astronomy and cosmology inference problems. Our results demonstrate that LtU-ILI yields competitive constraints on both synthetic and real science examples, indicating that these methods are ready out-of-the-box for application to new problems.
Looking ahead, this framework can be applied to accelerate the analyses of large surveys such as SDSS, Planck, VRO, Euclid, DESI, and JWST. We also aim to provide exhaustive benchmarks on the variety of problems described in Section 5, contributing to the formalization of machine learning analyses in astronomy and beyond. We intend to integrate LtU-ILI with state-of-the-art hyperparameter tuning (e.g., Akiba et al.(2019)Akiba, Sano, Yanase, Ohta, and Koyama) to provide an automatic, yet principled way for doing model selection. In addition, current approaches for calibrating posterior coverage rely on tweaking hyperparameters and retraining. An alternative approach is to implement recalibration methods (e.g., Masserano et al.(2022)Masserano, Dorigo, Izbicki, Kuusela, and Lee) which correctly calibrate posteriors given an independent validation set. We reserve implementation and analysis of these improvements to future work.
@@ Acknowledgements.
We thank Rosa Malandrino, Niall Jeffrey, and Justin Alsing for useful comments and suggestions on developing the codebase and writing the manuscript. This project was developed as part of the Simons Collaboration on “Learning the Universe.”

### v2

@@ 7 Conclusion
This manuscript is designed to serve as a reference document for future applications of LtU-ILI. In summary, we record detailed theoretical descriptions of every flavor of ILI (Section 2), describe the important functionalities of the codebase (Section 3), and provide guidance for experimental design amidst a broad landscape of estimators (Section 2.3) and backends (Section 3.4). We also have provided benchmarks and science examples on a variety of astronomy and cosmology inference problems. Our results demonstrate that LtU-ILI yields competitive constraints on both synthetic and real science examples, indicating that these methods are ready out-of-the-box for application to new problems.
Looking ahead, we aim to apply this framework to modern analyses of the problems described in Section 5, with an emphasis on observational applications to survey data from surveys such as SDSS, Planck, VRO, Euclid, DESI, and JWST. We aim to provide exhaustive benchmarks on these problems, contributing to the formalization of machine learning analyses in astronomy and beyond. We also intend to integrate LtU-ILI with state-of-the-art hyperparameter tuning (e.g., Akiba et al., 2019) to provide an automatic, yet principled way for doing model selection, as in Choustikov et al. (2024). This includes developing and implementing consolidated metrics for calibrating posterior coverage (e.g., Zhao et al., 2021; Lemos et al., 2024), to guarantee accurate posterior coverage through hyperparameter search. An alternative approach is to implement recalibration methods (e.g., Masserano et al., 2022) which correctly calibrate posteriors given an independent validation set.
We note that, while the current release of LtU-ILI is extensive, we hope to further improve the toolkit by incorporating additional Bayesian machine learning methods. For example, we do not currently include newer ILI procedures such as Truncated Marginal Neural Ratio Estimation (TMNRE; Miller et al., 2021, 2022) or score-based diffusion networks (e.g., Sharrock et al., 2022; Dax et al., 2023; Linhart et al., 2024), each of which has been shown to greatly improve upon the benchmarks that were investigated in Section 4. Also, during the development of this work, nbi (Zhang et al., 2023) was released as a neural posterior estimation toolkit focused on spectroscopy and light-curve analysis. Though we do not directly integrate nbi, we remark that the functionality of nbi is already encompassed and extended by LtU-ILI. Lastly, a major future addition to LtU-ILI will involve integration with implicit evidence estimators such as Evidence Networks (Jeffrey and Wandelt, 2024) and harmonic (McEwen et al., 2021; Spurio Mancini et al., 2023). With these features, users will be able to do Bayesian posterior inference and model comparison, all within one interface. We reserve implementation and analysis of each of these improvements to future work.
@@ Acknowledgements.
We thank Rosa Malandrino, Niall Jeffrey, and Justin Alsing for useful comments and suggestions on developing the codebase and writing the manuscript. This project was developed as part of the Simons Collaboration on “Learning the Universe.” The Flatiron Institute is supported by the Simons Foundation.